# Training Large Language Models to Reason In Parallel with Global Forking Tokens

### v1

@@ Abstract
Figure 1: An illustration of different supervised fine-tuning methods that aim to instill parallel reasoning capabilities from diverse reasoning traces. Compared to (1) standard SFT and (2) SFT with randomly assigned parallel thinking identifiers, (3) Set-Supervised Fine-Tuning leverages a self-supervised bipartite matching process to learn to initiate distinct reasoning modes conditioned on distinct <think i>, for each question. Self-supervised matching prevents collapse of distinct reasoning modes caused by ordering bias in a set of parallel reasoning traces.
@@ 1 Introduction
Large language models have recently improved reasoning by allocating more test-time compute to generate more tokens before producing the final answer (OpenAI, 2025). However, extended sequential scaling can lead to “overthinking”, where performance decreases beyond a certain sequence length (Ghosal et al., 2025; Chen et al., 2024a). To mitigate this, another scaling dimension based on repeated parallel sampling and aggregation (Wang et al., 2022; Brown et al., 2024) has shown success in further boosting reasoning performance. However, these methods rely on LLMs generating diverse yet correct solutions; as tasks become harder, a new mechanism is required to increase diversity. Recent work shows that only a minority of tokens in Chain-of-Thought reasoning (Wei et al., 2022) can act as forking tokens that lead to distinct reasoning modes (Wang et al., 2025), so as the problem becomes harder and the generation becomes longer, it can become substantially harder to sample them. Also, common practices to encourage diversity, typically through temperature scaling, inherently entail a diversity-accuracy trade-off, as the forking tokens that trigger diverse yet correct reasoning modes are typically located deeply within the sampling tree. Moreover, recent theoretical work also shows that increasing the temperature alone does not necessarily guarantee greater diversity unless the model is explicitly trained to ensure coverage. (Verine et al., 2025).
Building on these observations, we aim to leverage diverse reasoning traces to train for coverage (Guo et al., 2025; Google, 2025b). We introduce global forking tokens to initiate parallel reasoning traces and frame parallel reasoning as a set prediction problem. Specifically, given a question, an LLM, conditioned on a reserved set of tokens in a chosen ordering, generates $\mathrm{M}$ reasoning sequences in parallel, each aligned with one of $\mathrm{M}$ ground-truth reasoning traces. For each ordering, we compute the total autoregressive loss across the generated sequences. By enumerating all possible orderings, we identify the minimum loss, which defines the set language modeling loss, conditioned on the distinct forking tokens (Equation 5). This formulation naturally incorporates coverage into the training objective and is capable of learning global forking tokens that can serve as prompts to trigger reasoning modes that are both diverse and accurate. We operationalize this idea through our Set Supervised Fine-Tuning (SSFT) framework. Our main contributions are summarized below. We plan to open-source our code and model.
@@
  • We empirically demonstrate that, across multiple reasoning benchmarks, a model fine-tuned with SSFT outperforms a standard SFT model trained on the same distilled traces from diverse teacher models, improving $\mathrm{Pass@}1$, $\mathrm{Pass@}k$, and $\mathrm{Cons@}k$ (Section 3, Table 1 and Figure 4). We also find that naively fine-tuning using diverse reasoning traces without the set loss from bipartite matching can result in control tokens initiating the same collapsed reasoning mode (Figure 6). In contrast, the global forking tokens learned by SSFT reliably initiate distinct reasoning (Figure 5).
  • To facilitate training with variable-size parallel generation, we propose a scalable training implementation that does not increase VRAM usage. Instead of concatenating diverse reasoning traces, our algorithm can expand these variable-sized parallel generations along the batch dimension under distributed training (Appendix A.1).
@@ 2.1 Parallel Reasoning as Set of Next Token Prediction
Problem Setup. In this paper, our goal is not only to instill new reasoning capabilities into a model, but also to ensure that prompting with a set of reserved special tokens, in parallel with a question, elicits distinct reasoning traces. We call these global forking tokens ${\bm{g}}:=\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ instantiated as $\{\texttt{<think\,i>\,}\}_{i=1}^{\mathrm{N}}$ tags. We use ${\textnormal{g}}^{(i)}$ interchangeably with <thinki>, depending on context for clarity. We consider a setting with multiple sources of reasoning traces, obtained at low cost without human annotation by distilling from diverse teachers, sampling repeatedly, and potentially filtering with a verifiable metric such as correctness. We adopt this low-cost regime to highlight the effectiveness of our algorithm, though the method extends naturally to settings with well-annotated, human-labeled data. So our problem is to (1) learn to do a set of next-token predictions on multiple distinct yet correct reasoning traces ${\mathbf{R}}:=\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ in parallel for an input prompt $\mathbf{x}$}, and (2) ensure that distinct global forking tokens can uniquely initiate these distinct traces.
@@
To do this, we make a simple change to the NTP loss, which now has two requirements: (1) Permutation-invariance: It should not depend on the order of elements in $\mathbf{R}$} and $\bm{g}$}, so we don’t penalize a trace that incurs high NTP loss under one forking token if the model predicts it well when conditioned on another. (2)No shared global forking token: We want $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ to uniquely initiate distinct reasoning traces, so this requirement prevents conditioning on the same ${\textnormal{g}}^{(i)}$ when generating distinct traces given a question.
@@
To satisfy these requirements, we incorporate a subproblem in language modeling: finding the minimum cost bipartite matching configuration between the left vertices $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and the right vertices $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ where the cost of each edge between a left vertex $i$ and a right vertex $j$, is the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ and an input prompt $\mathbf{x}$}. A matching configuration is a set of edges connecting the left and right vertices where no two edges share a common vertex. Without loss of generality, we assume $\mathrm{N}\geq\mathrm{M}$ to simplify our notation. The total cost involves all vertices on the smaller side of the bipartite graph, so this allows us to write the summation from 1 to $\min\{\mathrm{N},\mathrm{M}\}$, which equals $\mathrm{M}$ under this assumption. We denote a matching configuration as a finite map $\bm{\sigma}:\{1,...,\mathrm{M}\}\rightarrow\{1,...,\mathrm{N}\}$ such that $\bm{\sigma}(j)=i\iff{\mathbf{r}}^{(j)}\text{ is paired with }{\textnormal{g}}^{(i)}$. Let $\bm{\mathfrak{S}}_{\mathrm{P}}:=\{\bm{\sigma}_{k}\}_{k=1}^{\mathrm{P}}$ denote all the $\binom{\mathrm{N}}{\mathrm{M}}\times\mathrm{M}!$ configurations of a bipartite graph. The total cost of each configuration represents the compatibility between $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ and $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ under this unique matching. Figure 2 shows an example of a matching configuration.
@@
Figure 2: An illustration of one SSFT training step. Step 1: We first construct the cost matrix by evaluating all pairwise combinations: for each ${\mathbf{r}}^{(j)}\in\{{\mathbf{r}}^{(1)},{\mathbf{r}}^{(2)},{\mathbf{r}}^{(3)},{\mathbf{r}}^{(4)}\}$ and each ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(1)},{\textnormal{g}}^{(2)},{\textnormal{g}}^{(3)},{\textnormal{g}}^{(4)},{\textnormal{g}}^{(5)},{\textnormal{g}}^{(6)}\}$, we compute the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ (Equation (4)). Then we use Hungarian algorithm to find $\hat{\bm{\sigma}}$ that minimizes the total bipartite matching cost. Here, this minimum is the sum of the losses highlighted in blue, which means $\hat{\bm{\sigma}}=\{({\textnormal{g}}^{(6)},{\mathbf{r}}^{(1)}),({\textnormal{g}}^{(5)},{\mathbf{r}}^{(2)}),({\textnormal{g}}^{(2)},{\mathbf{r}}^{(3)}),({\textnormal{g}}^{(3)},{\mathbf{r}}^{(4)})\}$. Step 2: We optimize $\bm{\theta}$} by backpropagating the set of NTP losses for ${\mathbf{r}}^{(j)}$, each conditioned on ${\textnormal{g}}^{(\hat{\bm{\sigma}}(j))}$.
@@ 2.2 SSFT: Minimizing Set of NTP Losses under Optimal Bipartite Matching
For the first step, we first compute all the entries in the cost matrix such as the one in Figure 2 and then apply the Hungarian algorithm (Kuhn, 1955) to efficiently find the optimal $\hat{\bm{\sigma}}$:
@@
After solving $\hat{\bm{\sigma}}$ for each $({\mathbf{x}},{\mathbf{R}})$, our second step optimizes model parameters $\bm{\theta}$} by backpropagating on the matching loss in Equation 5. The expectation is replaced by its sample mean over pairs of $({\mathbf{x}},\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}})$ in a mini-batch. In practice, we may use only the first $\mathrm{L}<{\mathrm{T}_{{\mathbf{r}}}}$ tokens in Equation 4 to compute the matching cost and find $\hat{\bm{\sigma}}$ when the training dataset is such that these early tokens reveal sufficient differences. However, we always optimize the matching loss for the full $\mathrm{T}_{{\mathbf{r}}}$} length.
@@
Remarks. The resulting model is not the same as a simple routing of the models independently trained with the nonoverlapping subsets of these traces. Firstly, SSFT allows positive transfer in representation learning within $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ even though they are matched to different $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$. Secondly, it is not optimal to distill reasoning traces from the same fixed sources for every question if the goal is to maximize both diversity and correctness. Our algorithm supports a variable number of target reasoning traces across training steps, with sources that may also change for each $\mathbf{x}$}. Thirdly, even if the two sets of traces, $\{{\mathbf{r}}^{(j)}_{a}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{a}$ and $\{{\mathbf{r}}^{(j)}_{b}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{b}$, are from the same sources, their optimal configurations $\hat{\bm{\sigma}}_{a}$ and $\hat{\bm{\sigma}}_{b}$ can still vary because a teacher model can reason differently under different questions. Lastly, we reserve more global forking tokens than the maximum number of traces ($\mathrm{N}>\mathrm{M}$), and empirically observe that all the forking tokens are being matched throughout the process. This is because the extra forking tokens can maximally intra-differentiate similar traces.
@@ 2.3 Inference with Learned Global Forking Tokens
Figure 3: (Left) $\bm{\mathfrak{S}}_{p}$ is the subset of bipartite matching configurations that are still computed optimal towards the end of training. These can be individually visualized as a bipartite (Middle) Learned matchings by SSFT-32B in Exp 3, obtained by aggregating all edges in $\bm{\mathfrak{S}}_{p}$. (Right) At test time, for $\mathrm{Pass@}1$, we prompt with ${\textnormal{g}}^{(i^{\star})}$ that has the most connected edges. For $\mathrm{Cons@}k$, we augment $i$th prompt by <think(i % $\mathrm{N}$)>.
@@ 3.1 Experiment Setup
Training Dataset. We use the 1,000 questions from s1k dataset Muennighoff et al. (2025). In addition to the R1 (Guo et al., 2025) and Gemini Flash (Google, 2024) traces provided by s1, we also use Claude Opus 4.0/4.1 (Anthropic, 2025) and GPT-OSS-120B (Agarwal et al., 2025) with high and medium reasoning effort to obtain a pool of distilled targets for the 1,000 questions. For each question, we generate two traces per source to populate the pool. We then sample four traces from this pool. We call this s1k-4mixed-reasoning dataset.
@@
Training Details. We fine-tune Qwen2.5-32B-Instruct (Yang et al., 2025a) for six epochs with a context length of 32,768. We reserve $\mathrm{N}=6$ global forking tokens and use $\mathrm{M}=4$ targets per question. To find the optimal bipartite matching for each input prompt, we consider only the first 1,000 tokens when computing the matching cost in Equation 4 for computational efficiency. We call this model SSFT-32B. We also include SSFT but choose a random bipartite matching at each step to fine-tune SSFT-32B (random $\bm{\sigma}$). Exact details on the pool of diverse distillation targets and selection procedure, as well as training hyperparameters, are provided in Appendix A.3.
@@
Baselines. All of our baselines use Qwen2.5-32B-Instruct as their base model, and only train on the 1k questions. Our baselines include two groups: (Single-Target $\triangle$) models trained with one trace per question and (Multi-Target $\bigstar$) models trained with four traces per question. For (Single-Target), we include s1.1-32B (Muennighoff et al., 2025), which uses 1k DeepSeek-R1 traces. We also fine-tuned an SFT-OSS-distill-32B baseline that trains only on the 1k GPT-OSS traces with high reasoning effort, as these traces achieved the highest correctness on the 1k questions based on an evaluation by Claude 3.5 Sonnet comparing each attempt against the reference answer. For (Multi-Target), we use our s1k-4mixed-reasoning to fine-tune SFT-mixed-distill-32B using standard SFT with one <think> token, duplicating each question and treating the four traces as four individual data points. We also include Multiverse-32B (Yang et al., 2025b), which prompts Gemini 2.5-Pro (Google, 2025a) to transform 1k sequential CoTs into parallel CoTs as their training data.
@@
Evaluation Setup. Our evaluation tasks consist of AIME24/AIME25 (Ye et al., 2025), MATH-500 (Hendrycks et al., 2021), and GPQA-Diamond (Rein et al., 2024). We use LightEval (Habib et al., 2023) as our evaluation framework with generation configurations: temperature=0.7 used in (Guha et al., 2025), top_p=0.95, max length=32768. For $\mathrm{Pass@}1$ accuracy without any parallel test-time compute, we select learned ${\textnormal{g}}^{(1)}$ for SSFT-32B and ${\textnormal{g}}^{(4)}$ for SSFT-32B (random $\bm{\sigma}$) based on Equation 6. For each $\mathrm{Pass@}1$ accuracy, we compute the average performance over 32 generations. For $\mathrm{Cons@}6$, which applies each of the six global forking tokens once in our method and uses six generations for the baselines, we compute the average over 11 sets of generations to reduce variance in the results. We refer to this as Pass@1 of Native Cons@6 using a similar terminology as the concurrent work (Wen et al., 2025). Appendix  A.4 presents an example of the parallel generations.

### v2

@@ Abstract
Figure 1: An illustration of different supervised fine-tuning methods that aim to instill parallel reasoning capabilities from diverse reasoning traces for the same question. Compared to (1) standard SFT and (2) SFT with randomly assigned parallel thinking identifiers, (3) Set-Supervised Fine-Tuning leverages a self-supervised bipartite matching process to learn to maximally differentiate the reasoning modes conditioned on distinct <think i>, for each question. Self-supervised matching prevents collapse of maximally distinct reasoning modes caused by ordering bias from randomly or manually assigning a reasoning mode identifier to a reasoning trace. The learned global forking tokens are considered emergent because we observe that similar reasoning modes over the traces for different questions cluster to match the same <think,i> at convergence, even though we do not manually identify such modes or incorporate any regularization term to maintain consistency in matching over different input prompts.
@@ 1 Introduction
Large language models have recently improved reasoning by allocating more test-time compute to generate more tokens before producing the final answer (OpenAI2025mini). However, extended sequential scaling can lead to “overthinking”, where performance decreases beyond a certain sequence length (ghosal2025doesoverthinkinghelp; chen2024notoverthinkthatmuch). To mitigate this, another scaling dimension based on repeated parallel sampling and aggregation (wang2022selfconsistency; brown2024large_repeatedsampling) has shown success in further boosting reasoning performance. However, these methods rely on LLMs generating diverse yet correct solutions; as tasks become harder, a new mechanism is required to increase diversity. Recent work shows that only a minority of tokens in Chain-of-Thought reasoning (wei2022chain) can act as forking tokens that lead to distinct reasoning modes (wang2025beyond_80_20_rule; wang2025emergenthierarchicalreasoning), so as the problem becomes harder and the generation becomes longer, it can become substantially harder to sample them. Also, common practices to encourage diversity, typically through temperature scaling, inherently entail a diversity-accuracy trade-off, as the forking tokens that trigger diverse yet correct reasoning modes are typically located deeply within the sampling tree. Moreover, recent theoretical work also shows that increasing the temperature alone does not necessarily guarantee greater diversity unless the model is explicitly trained for coverage. (verine2025improvingwhentemperaturefailschangetheloss).
Building on these observations, we aim to leverage diverse reasoning traces to train for coverage (guo2025deepseekR1; gemini25). We introduce global forking tokens to initiate parallel reasoning traces and frame parallel reasoning as a set prediction problem. Specifically, given a question, an LLM, conditioned on a reserved set of tokens in a chosen ordering, generates $\mathrm{M}$ reasoning sequences in parallel, each aligned with one of $\mathrm{M}$ ground-truth reasoning traces. For each ordering, we compute the total autoregressive loss across the generated sequences. By enumerating all possible orderings, we identify the minimum loss, which defines the set language modeling loss, conditioned on the distinct forking tokens (Equation 5). This formulation naturally incorporates coverage into the training objective and is capable of learning global forking tokens that can serve as prompts to trigger reasoning modes that are both diverse and accurate. We operationalize this idea through our Set Supervised Fine-Tuning (SSFT) framework. Our main contributions are summarized below.
@@
  • We empirically demonstrate that, across multiple reasoning benchmarks, a model fine-tuned with SSFT outperforms a standard SFT model trained on the same distilled traces from diverse teacher models, improving $\mathrm{Pass@}1$, $\mathrm{Pass@}k$, and $\mathrm{Cons@}k$ (Section 3, Table 1 and Figure 4). We also find that naively fine-tuning using diverse reasoning traces without the set loss from bipartite matching can result in such control tokens initiating the same collapsed reasoning mode (Figure 6). In contrast, the global forking tokens learned by SSFT maximally differentiate reasoning modes (Figure 5).
  • To facilitate training where different input prompts can use different numbers of parallel reasoning traces, which can cause difficulty in using gradient accumulation under distributed training, we also include a simple implementation based on the HuggingFace repository. Our implementation does not concatenate a variable number of reasoning traces along the sequence dimension and modify causal attention, which would require sequence parallelism and more complicated fused kernels to avoid out-of-memory for training with long reasoning traces (Appendix A.1).
@@ 2.1 Parallel Reasoning as Set of Next Token Prediction
Problem Setup. In this paper, our goal is not only to instill new reasoning capabilities into a model, but also to ensure that prompting with a set of reserved special tokens, in parallel with a question, elicits distinct reasoning traces. We call these global forking tokens ${\bm{g}}:=\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ instantiated as $\{\texttt{<think\,i>\,}\}_{i=1}^{\mathrm{N}}$ tags. We use ${\textnormal{g}}^{(i)}$ interchangeably with <thinki>, depending on context for clarity. We consider a setting with multiple sources of reasoning traces, obtained at low cost without human annotation by distilling from diverse teachers, sampling repeatedly, and potentially using rejection sampling to filter out low-quality or incorrect traces. We adopt this low-cost regime to highlight the effectiveness of our algorithm, though the method extends naturally to settings with well-annotated, human-labeled data. So our problem is to (1) learn to do a set of next-token predictions on multiple distinct yet correct reasoning traces ${\mathbf{R}}:=\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ in parallel for an input prompt $\mathbf{x}$}, and (2) ensure that distinct global forking tokens can uniquely initiate these distinct traces.
@@
To do this, we make a simple change to the NTP loss, which now has two requirements: (1) Permutation-invariance: It should not depend on the order of elements in $\mathbf{R}$} and $\bm{g}$}, so we don’t penalize a trace that incurs high NTP loss under one global forking token if the model predicts it well when conditioned on another. (2)No shared global forking token: We want $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ to uniquely initiate distinct reasoning traces, so this requirement prevents conditioning on the same ${\textnormal{g}}^{(i)}$ when generating distinct traces given a question.
@@
To satisfy these requirements, we formally incorporate a subproblem in language modeling: finding the minimum cost bipartite matching configuration between the left vertices $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ and the right vertices $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ where the cost of each edge between a left vertex $i$ and a right vertex $j$, is the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ and an input prompt $\mathbf{x}$}. A matching configuration is a set of edges connecting the left and right vertices where no two edges share a common vertex. Without loss of generality, we assume $\mathrm{N}\geq\mathrm{M}$ to simplify our notation. The total cost involves all vertices on the smaller side of the bipartite graph, so this allows us to write the summation from 1 to $\min\{\mathrm{N},\mathrm{M}\}$, which equals $\mathrm{M}$ under this assumption. We denote a matching configuration as a finite map $\bm{\sigma}:\{1,...,\mathrm{M}\}\rightarrow\{1,...,\mathrm{N}\}$ such that $\bm{\sigma}(j)=i\iff{\mathbf{r}}^{(j)}\text{ is paired with }{\textnormal{g}}^{(i)}$. Let $\bm{\mathfrak{S}}_{\mathrm{P}}:=\{\bm{\sigma}_{k}\}_{k=1}^{\mathrm{P}}$ denote all the $\binom{\mathrm{N}}{\mathrm{M}}\times\mathrm{M}!$ configurations of a bipartite graph. The total cost of each configuration represents the compatibility between $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ and $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$ under this unique matching. Figure 2 shows an example of a matching configuration.
@@
Figure 2: An illustration of one SSFT training step. Step 1: We first construct the cost matrix by evaluating all pairwise combinations: for each ${\mathbf{r}}^{(j)}\in\{{\mathbf{r}}^{(1)},{\mathbf{r}}^{(2)},{\mathbf{r}}^{(3)},{\mathbf{r}}^{(4)}\}$ and each ${\textnormal{g}}^{(i)}\in\{{\textnormal{g}}^{(1)},{\textnormal{g}}^{(2)},{\textnormal{g}}^{(3)},{\textnormal{g}}^{(4)},{\textnormal{g}}^{(5)},{\textnormal{g}}^{(6)}\}$, we compute the NTP loss of ${\mathbf{r}}^{(j)}$ conditioned on ${\textnormal{g}}^{(i)}$ (Equation (4)). Then we use Hungarian algorithm to find $\hat{\bm{\sigma}}$ that minimizes the total bipartite matching cost. Here, this minimum is the sum of the losses highlighted in blue, which means $\hat{\bm{\sigma}}=\{({\textnormal{g}}^{(6)},{\mathbf{r}}^{(1)}),({\textnormal{g}}^{(5)},{\mathbf{r}}^{(2)}),({\textnormal{g}}^{(2)},{\mathbf{r}}^{(3)}),({\textnormal{g}}^{(3)},{\mathbf{r}}^{(4)})\}$. Step 2: We optimize $\bm{\theta}$} by backpropagating the set of NTP losses for ${\mathbf{r}}^{(j)}$, each conditioned on ${\textnormal{g}}^{(\hat{\bm{\sigma}}(j))}$. This is the Hungarian loss in Equation 5.
@@ 2.2 SSFT: Minimizing Set of NTP Losses under Optimal Bipartite Matching
For the first step, we compute all the entries in the cost matrix such as the one in Figure 2 and then apply the Hungarian algorithm (hungarian) to efficiently find the optimal $\hat{\bm{\sigma}}$:
@@
After solving $\hat{\bm{\sigma}}$ for each $({\mathbf{x}},{\mathbf{R}})$, our second step optimizes model parameters $\bm{\theta}$} by backpropagating the Hungarian loss in Equation 5. In Algorithm 1, the expectation is replaced by its sample mean over pairs of $({\mathbf{x}},\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}})$ in a mini-batch. In practice, we may use only the first $\mathrm{L}<{\mathrm{T}_{{\mathbf{r}}}}$ tokens in Equation 4 to compute the matching cost and find $\hat{\bm{\sigma}}$ when the training dataset is such that these early tokens reveal sufficient differences. However, we always optimize the Hungarian loss for the full $\mathrm{T}_{{\mathbf{r}}}$} length.
@@
Remarks. The resulting model is not the same as a simple routing of the models independently trained with the nonoverlapping subsets of these traces. Firstly, SSFT allows positive transfer in representation learning within $\{{\mathbf{r}}^{(j)}\}_{j=1}^{\mathrm{M}}$ even though they are matched to different $\{{\textnormal{g}}^{(i)}\}_{i=1}^{\mathrm{N}}$. Secondly, it is not optimal to distill reasoning traces from the same fixed sources for every question if the goal is to maximize both diversity and correctness. Some traces can be filtered by rejection sampling. Our algorithm supports a variable number of target reasoning traces across training steps, with sources that may also change for each $\mathbf{x}$}. Thirdly, even if the two sets of traces, $\{{\mathbf{r}}^{(j)}_{a}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{a}$ and $\{{\mathbf{r}}^{(j)}_{b}\}_{j=1}^{\mathrm{M}}$ for ${\mathbf{x}}_{b}$, are from the same sources, their optimal configurations $\hat{\bm{\sigma}}_{a}$ and $\hat{\bm{\sigma}}_{b}$ can still vary because a teacher model can reason differently under different questions. Lastly, we reserve more global forking tokens than the maximum number of traces ($\mathrm{N}>\mathrm{M}$), and empirically observe that all the forking tokens are being matched throughout the process. This is because the extra forking tokens can maximally intra-differentiate similar traces.
@@ 2.3 Inference with Learned Global Forking Tokens
Figure 3: (Left) $\bm{\mathfrak{S}}_{p}$ is the subset of bipartite matching configurations that are still computed optimal towards the end of training. These can be individually visualized as a bipartite (Middle) Learned matchings by SSFT-32B in Exp 3, obtained by aggregating all edges in $\bm{\mathfrak{S}}_{p}$. (Right) At test time, for $\mathrm{Pass@}1$, we prompt with ${\textnormal{g}}^{(i^{\star})}$ that has the most connected edges. For $\mathrm{Cons@}k$, we prompt $i$-th parallel generation with <think(i % $\mathrm{N}$)>.
@@ 3.1 Experiment Setup
Training Dataset. We use the 1,000 questions from s1k dataset muennighoff2025s1_simpletesttime. In addition to the R1 (guo2025deepseekR1) and Gemini Flash (geminithinking) traces provided by s1, we also use Claude Opus 4.0/4.1 (claude_41_opus) and GPT-OSS-120B (agarwal2025gptoss) with high and medium reasoning effort to obtain a pool of distilled targets for the 1,000 questions. For each question, we generate two traces per source to populate the pool. We then sample four traces from this pool. We call this s1k-4mixed-reasoning dataset.
@@
Training Details. We fine-tune Qwen2.5-32B-Instruct (yang2025qwen3) for six epochs with a context length of 32,768. We reserve $\mathrm{N}=6$ global forking tokens and use $\mathrm{M}=4$ targets per question. To find the optimal bipartite matching for each input prompt, we consider only the first 1,000 tokens when computing the matching cost in Equation 4 for computational efficiency. We call this model SSFT-32B. We also include SSFT but choose a random bipartite matching at each step to fine-tune SSFT-32B (random $\bm{\sigma}$). Exact details on the pool of diverse distillation targets and selection procedure, as well as training hyperparameters, are provided in Appendix A.3.
@@
Baselines. All of our baselines use Qwen2.5-32B-Instruct as their base model, and only train on the 1k questions. Our baselines include two groups: (Single-Target $\triangle$) models trained with one trace per question and (Multi-Target $\bigstar$) models trained with four traces per question. For (Single-Target), we include s1.1-32B (muennighoff2025s1_simpletesttime), which uses 1k DeepSeek-R1 traces. We also fine-tuned an SFT-OSS-distill-32B baseline that trains only on the 1k GPT-OSS traces with high reasoning effort, as these traces achieved the highest correctness on the 1k questions based on an evaluation by Claude 3.5 Sonnet comparing each attempt against the reference answer. For (Multi-Target), we use our s1k-4mixed-reasoning to fine-tune SFT-mixed-distill-32B using standard SFT with one <think> token, duplicating each question and treating the four traces as four individual data points. We also include Multiverse-32B (yang2025multiverse), which prompts Gemini 2.5-Pro (Google:Gemini2.5Pro2025) to transform 1k sequential CoTs into parallel CoTs as their training data.
@@
Evaluation Setup. Our evaluation tasks consist of AIME24/AIME25 (ye2025aimepreview), MATH-500 (hendrycks2021measuringmath500), and GPQA-Diamond (rein2024gpqa). We use LightEval (lighteval) as our evaluation framework with generation configurations: temperature=0.7 used in (guha2025openthoughts), top_p=0.95, max length=32768. For $\mathrm{Pass@}1$ accuracy without any parallel test-time compute, we select learned ${\textnormal{g}}^{(1)}$ for SSFT-32B and ${\textnormal{g}}^{(4)}$ for SSFT-32B (random $\bm{\sigma}$) based on Equation 6. For each $\mathrm{Pass@}1$ accuracy, we compute the average performance over 32 generations. For $\mathrm{Cons@}6$, which applies each of the six global forking tokens once in our method and uses six generations for the baselines, we compute the average over 11 sets of generations to reduce variance in the results. We refer to this as Average of Native Cons@6, which corresponds to the Pass@1 metric with parallel generations in concurrent work (wen2025parathinker). Appendix  A.4 presents an example of the parallel generations by SSFT-32B.